# TS-RAG: Retrieval-Augmented Generation based Time Series Foundation Models are Stronger Zero-Shot Forecaster

### v3

@@ Abstract
Large Language Models (LLMs) and Foundation Models (FMs) have recently become prevalent for time series forecasting tasks. While fine-tuning LLMs enables domain adaptation, they often struggle to generalize across diverse and unseen datasets. Moreover, existing Time Series Foundation Models (TSFMs) still face challenges in handling non-stationary dynamics and distribution shifts, largely due to the lack of effective mechanisms for adaptation. To this end, we present TS-RAG, a retrieval-augmented generation framework for time series forecasting that enhances the generalization and interpretability of TSFMs. Specifically, TS-RAG leverages pre-trained time series encoders to retrieve semantically relevant segments from a dedicated knowledge base, enriching the contextual representation of the input query. Furthermore, we propose an Adaptive Retrieval Mixer (ARM) module that dynamically fuses the retrieved patterns with the TSFM’s internal representation, improving forecasting accuracy without requiring task-specific fine-tuning. Thorough empirical studies on seven public benchmark datasets demonstrate that TS-RAG achieves state-of-the-art zero-shot forecasting performance, outperforming the existing TSFMs by up to 6.84% across diverse domains while also providing desirable interpretability.
@@ 1 Introduction
Time series forecasting, which aims to predict future values of a sequence based on its past observations, plays a critical role in various real-world applications, e.g., finance, healthcare, energy management, and climate science. Modeling time series data essentially captures the temporal dependency patterns in the form of trend, seasonality, autocorrelation, etc, to make accurate predictions and generalize across different datasets. In the past, a substantial amount of effort has been made to tackle this problem. Traditional statistical methods such as AutoRegressive Integrated Moving Average (ARIMA) [1] work well for stationary time series but struggle with complex dependencies and non-linear patterns. Machine learning approaches, such as Random Forest [2] and XGBoost [3], can handle external covariates of features but fail to capture long-range dependencies. Deep learning techniques, including Long Short-Term Memory (LSTM) [4], Gated Recurrent Units (GRUs) [5], Temporal Convolutional Networks (TCNs) [6], Graph Neural Networks (GNN) based models [7, 8], and transformer based models [9, 10] are typically trained within specific domains and often struggle to generalize effectively to diverse, unseen datasets. More recently, there is a prevalent interest in adapting Large Language Models (LLMs) [11, 12, 13] for time series tasks [14, 15] and developing Foundation Models (FM) tailored for time series data [16, 17, 18]. Although both LLMs and FM have shown great promise in improving forecasting accuracy and handling complex temporal dynamics, they still face immense barriers when applied to zero-shot forecasting tasks, limiting their real-world applicability.
@@
Specifically, recent work [19] has explored the potential usage of LLMs as zero-shot forecasters for time series tasks. While fine-tuning LLMs can facilitate adaptation and yield better performance over specific datasets, such methods often struggle to generalize across diverse, unseen domains [20, 21, 22], and typically involve substantial computational overhead, even when applied to limited data. Time Series Foundation Models (TSFMs)  [23, 24, 25, 26] have emerged as a promising alternative by learning time series representations across a wide range of time series data. However, they lack inherent mechanisms for domain adaptation, as they cannot incorporate external contextual knowledge dynamically. This shortcoming limits their capability in handling complex, non-stationary, and evolving time series patterns. Furthermore, most TSFMs offer limited interpretability, which poses challenges for their deployment in high-stakes or decision-critical applications.
@@
Recently, Retrieval-Augmented Generation (RAG) [27] has demonstrated significant success across various Natural Language Processing (NLP) tasks [28, 29, 30, 31]. By retrieving relevant document segments and incorporating them into prompts, RAG refines the existing prompts and enables LLMs to generate more informed, context-aware outputs, thereby improving both accuracy and adaptability in diverse applications. Motivated by this paradigm, we propose TS-RAG, a retrieval-augmented generation based time series forecasting framework. TS-RAG dynamically retrieves semantically relevant time series patterns and integrates them into the forecasting pipeline, enabling strong zero-shot performance without the need for fine-tuning. In addition, it significantly enhances the interpretability of TSFMs by providing contextual evidence for their predictions.
@@ 2 Related Work
Time Series Foundation Models Existing LLM-based time series forecasters [20, 32, 33, 21, 34, 22] have demonstrated remarkable achievements in in-domain time series analysis. However, the domain adaptation challenges and significant computational costs associated with LLM-based models have motivated the emergence of time series foundation models as a more efficient and scalable alternative. Inspired by recent advancements in Natural Language Processing (NLP) and Vision Transformers [35], Time Series Foundation Models have rapidly developed and drawn significant attention. These models have demonstrated strong generalization capabilities across diverse datasets, leading to substantial progress in time series forecasting. Lag-Llama [16] and TimeGPT-1 [17] are pioneering forecasting foundation models, pre-trained on extensive time series datasets spanning multiple domains. Lag-Llama utilizes lagged time series features and the LLaMA architecture [36], while TimeGPT-1 adopts an encoder-decoder transformer structure to handle forecasting tasks effectively. Following the paradigm of patching tokenization and optimization, models such as TimesFM [23], MOMENT [37], Timer [38, 39], and Sundial [40] first patch and embed continuous time series values, and subsequently model the output distributions and perform point forecasting. To further improve efficiency, Tiny Time Mixers (TTMs)[24] train a compact foundation model, while TimeMOE[41] adopts a sparse mixture-of-experts architecture that activates only a subset of expert networks during inference, thereby maintaining strong performance with reduced computational cost. However, deterministic predictions usually cannot satisfy the requirement of decision-making. To address this limitation, Moirai [26] trains a probabilistic model that captures a mixture of distributions. Building on language modeling techniques, Chronos [25] discretizes time series through scaling and quantization, and models the resulting categorical distributions using cross-entropy loss. To enhance inference efficiency and forecasting accuracy, Chronos-bolt replaces discrete tokenization with a patch-based input strategy and utilizes decoder representations to produce quantile forecasts over multiple future steps, achieving improved performance over its predecessor. These methods, however, lack inherent mechanisms to incorporate external contextual knowledge dynamically to facilitate zero-shot learning and suffer from limited interpretability.
@@
Retrieval-Augmented for Time Series Forecasting Although both LLMs and TSFMs have achieved strong performance in time series forecasting, they can still struggle in scenarios involving non-stationarity or distribution shifts. To address these challenges, Retrieval-Augmented Generation (RAG) techniques [27] offer a promising approach to incorporate external knowledge and enhance generalization and robustness. Recent works have explored the integration of retrieval mechanisms into time series forecasting. Among them, several approaches rely on fine-tuning to adapt the model to downstream tasks, including ReTime [42], RATD [43], TimeRAG [44], and RAFT [45]. ReTime [42] proposes relational retrieval and content synthesis for retrieval-based time series forecasting. RATD [43] utilizes retrieved historical time series to guide the denoising process of diffusion models, enhancing forecasting performance. RAFT [45] integrates retrieval with a multi-resolution forecasting framework. These methods typically construct the retrieval database from the training set of the target task and require fine-tuning for model adaptation. Similarly, TimeRAG [44] incorporates retrieved sequences into the forecasting process by using a frozen LLM backbone, and introduces a trainable reprogramming layer to align time series and text modalities.
@@
RAF [46] is a pioneering work that introduces a retrieval-augmented framework for zero-shot time series forecasting. It utilizes Chronos as the backbone model and constructs the augmented input by directly concatenating the processed retrieved context with the original time series input. However, this approach may face scalability and efficiency challenges when applied to large-scale retrieval databases. Moreover, RAG techniques for time series forecasting remain underexplored, particularly in terms of knowledge base construction, augmentation strategies, and their integration with TSFMs to facilitate zero-shot forecasting. To bridge these gaps, we propose TS-RAG, a retrieval-augmented framework specifically designed to enhance zero-shot forecasting capabilities of TSFMs.
@@ 3.1 Construction of the Retrieval Knowledge Base
TSFMs are typically pretrained on a multi-domain time series dataset to enhance their generalization capability. For TS-RAG, we adopt a similar strategy, i.e., construct a multi-domain dataset and specifically focus on learning the ARM augmentation module. We leverage the pretraining dataset of Chronos [25], which utilizes TSMixup to randomly combine time series data points from various domains. This approach enhances data diversity by blending different patterns, thereby improving the model’s ability to generalize. Given the fact that the trainable parameters of TS-RAG are significantly less than those in the TSFM backbone and the retrieval encoder (more discussions are available in Appendix B.5), we uniformly sample a subset from the Chronos pretraining dataset to serve as the pretraining dataset for TS-RAG. Based on this subset, we can further draw a subset to construct the retrieval knowledge database for TS-RAG, which will be used in the inference (forecasting) stage.
@@ 3.2 Architecture of TS-RAG Framework
Within TS-RAG, a TSFM has three key components [47]: an encoding layer, which may include normalization and embedding layers to preprocess and transform the input time series; a backbone, typically implemented as a transformer based model (i.e. GPT [48], T5 [49], Llama [12] etc.) to extract temporal representations; and a projection layer, often implemented as a multi-layer perceptron (MLP), which maps the temporal representations from the backbone to the final prediction values.
@@ 4.1 Experimental Setup
Datasets and Retrieval Knowledge Base. For the pretraining dataset, we first uniformly sample 50 million data points from the Chronos pretraining dataset [25] and further uniformly sample a subset of 5 million data points to construct the multi-domain retrieval knowledge base. To facilitate efficient indexing and retrieval, both the pretraining dataset and the retrieval knowledge base are segmented using a predefined context window. This process results in a total of 26 million pretraining data pairs and 2.8 million retrieval knowledge base pairs.
@@
Baselines. In practice, we use Chronos-Bolt, one of the state-of-the-art TSFMs, as the backbone of TS-RAG, as it achieves competitive performance in our evaluations. TS-RAG is designed to be compatible with any general TSFM. While our main experiments primarily use Chronos-Bolt as the backbone due to its strong empirical performance, we also verify TS-RAG’s effectiveness with MOMENT [37] in Appendix B.1. For comparison, we also report the zero-shot performance of other TSFMs, including TTM [24], TimesFM [23], Moirai [26], Chronos [25], Chronos-Bolt [25], MOMENT [37], and Time-MoE [41].
@@ 4.3 Ablation Studies
We further conduct additional ablation studies to investigate the impact of retriever configurations, including the number of retrieved sequences, retrieval lookback length, and retriever encoder choice. The results confirm that while TS-RAG benefits from longer lookback windows and a moderate number of retrieved sequences, it is robust to the choice of retriever encoder. Detailed results and analysis are provided in Appendices B.3, B.4, and B.5.
@@ 4.4 Comparison with RAF
To further evaluate the performance of TS-RAG, we compare it with RAF (Retrieval Augmented Forecasting) [46], a recent retrieval-augmented method for zero-shot time series forecasting. We conduct the comparison from two aspects: effectiveness and efficiency. More implementation details can be found in Appendix A.3.
@@ A.1 TS-RAG Implementation Details
[…]
@@ A.2 TSFM Baseline Introduction
  • Chronos-bolt [25]: Chronos-bolt is a subsequent version of Chronos, which can handle patch-based and uses decoder representations to generate quantile forecasts across multiple future steps, improving forecast accuracy over Chronos.
  • MOMENT [37]: MOMENT uses a masking modeling technique for zero-shot forecasting by appending a lookback series with a mask that matches the length of the forecast. It involves pretraining a Transformer encoder model univariately on the "Time Series Pile" datasets, which include a wide variety of time series data.
  • TTM [24]: TTM pre-trains a compact model based on the lightweight TSMixer architecture. It incorporates adaptive patching, diverse resolution sampling, and resolution prefix tuning to pretrain successfully on a small dataset.
  • Moirai [26]: Moirai pretrains the Transformer encoder on the “LOTSA” dataset, which includes 27B time points, by masking the forecast horizon of each target channel and performing mask reconstruction.
  • TimesFM [23]: TimesFM employs a decoder-style attention model and is pre-trained in a univariate manner on a large group of both real-world and synthetic datasets.
  • Chronos [25]: Chronos is a probabilistic time series foundation model.Chronos tokenizes the input time series in a quantized manner and processes these tokens using the T5 model [49]. Chronos is trained on an extensive corpus of collected and synthetic time series data and has great generalization ability.
  • Time-MOE [41]: Time-MoE consists of a family of decoder-only time series foundation models with a mixture-of-experts architecture, designed to operate in an auto-regressive manner, enabling universal forecasting with arbitrary prediction horizons and long context lengths.

### v4

@@ Abstract
Large Language Models (LLMs) and Foundation Models (FMs) have recently become prevalent for time series forecasting tasks. While fine-tuning LLMs enables domain adaptation, they often struggle to generalize across diverse and unseen datasets. Moreover, existing Time Series Foundation Models (TSFMs) still face challenges in handling non-stationary dynamics and distribution shifts, largely due to the lack of effective mechanisms for adaptation. To this end, we present TS-RAG, a retrieval-augmented generation framework for time series forecasting that enhances the generalization and interpretability of TSFMs. Specifically, TS-RAG leverages pre-trained time series encoders to retrieve semantically relevant segments from a dedicated knowledge base, enriching the contextual representation of the input query. Furthermore, we propose an Adaptive Retrieval Mixer (ARM) module that dynamically fuses the retrieved patterns with the TSFM’s internal representation, improving forecasting accuracy without requiring task-specific fine-tuning. Thorough empirical studies on seven public benchmark datasets demonstrate that TS-RAG achieves state-of-the-art zero-shot forecasting performance, outperforming the existing TSFMs by up to 6.84% across diverse domains while also providing desirable interpretability. Our code and data are available at: https://github.com/UConn-DSIS/TS-RAG.
@@ 1 Introduction
Time series forecasting, which aims to predict future values of a sequence based on its past observations, plays a critical role in various real-world applications, e.g., finance rezaei2021stock, healthcare jin2018treatment, energy management tzelepi2023deeplearningenergytimeseries, and climate science sun2022accurate. Modeling time series data essentially captures the temporal dependency patterns in the form of trend, seasonality, autocorrelation, etc, to make accurate predictions and generalize across different datasets. In the past, a substantial amount of effort has been made to tackle this problem. Traditional statistical methods such as AutoRegressive Integrated Moving Average (ARIMA) Whittle1951 work well for stationary time series but struggle with complex dependencies and non-linear patterns. Machine learning approaches, such as Random Forest breiman2001random and XGBoost chen2016xgboost, can handle external covariates of features but fail to capture long-range dependencies. Deep learning techniques, including Long Short-Term Memory (LSTM) hochreiter1997long, Gated Recurrent Units (GRUs) cho2014learning, Temporal Convolutional Networks (TCNs) lea2016temporal, Graph Neural Networks (GNN) based models cao2020spectral; shang2021discrete, and transformer based models vaswani2017attention; zhou2021informer are typically trained within specific domains and often struggle to generalize effectively to diverse, unseen datasets. More recently, there is a prevalent interest in adapting Large Language Models (LLMs) brown2020language; touvron2023llama2; Achiam2023GPT4TR; sun2025adapting for time series tasks (gruver2024large; jiang2024empowering) and developing Foundation Models (FM) tailored for time series data (rasul2023lag; garza2023timegpt; liutimer; liang_foundation_survey_2024). Although both LLMs and FM have shown great promise in improving forecasting accuracy and handling complex temporal dynamics, they still face immense barriers when applied to zero-shot forecasting tasks, limiting their real-world applicability.
Figure 1: Overview of the proposed TS-RAG framework. Given an input time series as the query, the retriever accesses a knowledge base to obtain semantically related information. The retrieved knowledge is subsequently integrated into a frozen time series foundation model (which can adopt various architectures and design choices) to enhance forecasting performance.
@@
Specifically, recent work gruver2024largelanguagemodelszeroshot has explored the potential usage of LLMs as zero-shot forecasters for time series tasks. While fine-tuning LLMs can facilitate adaptation and yield better performance over specific datasets, such methods often struggle to generalize across diverse, unseen domains zhou2023one; jin2023time; pan2024s, and typically involve substantial computational overhead, even when applied to limited data. Time Series Foundation Models (TSFMs)  das2023decoder; ekambaram2024tiny; ansari2024chronos; woo2024unified have emerged as a promising alternative by learning time series representations across a wide range of time series data. However, they lack inherent mechanisms for domain adaptation, as they cannot incorporate external contextual knowledge dynamically. This shortcoming limits their capability in handling complex, non-stationary, and evolving time series patterns. Furthermore, most TSFMs offer limited interpretability, which poses challenges for their deployment in high-stakes or decision-critical applications.
@@
Recently, Retrieval-Augmented Generation (RAG) lewis2020retrieval has demonstrated significant success across various Natural Language Processing (NLP) tasks RAG_LLM_2024; LLM_Know_RAG_2024; he2024gretriever; FT_or_RAG_2024. By retrieving relevant document segments and incorporating them into prompts, RAG refines the existing prompts and enables LLMs to generate more informed, context-aware outputs, thereby improving both accuracy and adaptability in diverse applications. Motivated by this paradigm, we propose TS-RAG, a retrieval-augmented generation based time series forecasting framework. TS-RAG dynamically retrieves semantically relevant time series patterns and integrates them into the forecasting pipeline, enabling strong zero-shot performance without the need for fine-tuning. In addition, it significantly enhances the interpretability of TSFMs by providing contextual evidence for their predictions. An overview of the proposed framework is illustrated in Figure 1, where a retriever queries a knowledge base for semantically related time-series information, and the retrieved knowledge is used to augment a frozen TSFM that can adopt diverse architectural designs.
@@ 2 Related Work
Time Series Foundation Models Existing LLM-based time series forecasters zhou2023one; xue2023promptcast; chang2023llm4ts; jin2023time; sun2023test; pan2024s; niulangtime have demonstrated remarkable achievements in in-domain time series analysis. However, the domain adaptation challenges and significant computational costs associated with LLM-based models have motivated the emergence of time series foundation models as a more efficient and scalable alternative. Inspired by recent advancements in Natural Language Processing (NLP) and Vision Transformers chen2025visionts, Time Series Foundation Models have rapidly developed and drawn significant attention. These models have demonstrated strong generalization capabilities across diverse datasets, leading to substantial progress in time series forecasting. Lag-Llama (rasul2023lag) and TimeGPT-1 (garza2023timegpt) are pioneering forecasting foundation models, pre-trained on extensive time series datasets spanning multiple domains. Lag-Llama utilizes lagged time series features and the LLaMA architecture (touvron2023llama), while TimeGPT-1 adopts an encoder-decoder transformer structure to handle forecasting tasks effectively. Following the paradigm of patching tokenization and optimization, models such as TimesFM (das2023decoder), MOMENT (goswami2024moment), Timer (Timer; Timer-XL), and Sundial (liu2025sundial) first patch and embed continuous time series values, and subsequently model the output distributions and perform point forecasting. To further improve efficiency, Tiny Time Mixers (TTMs)(ekambaram2024tiny) train a compact foundation model, while TimeMOE(shi2024timemoe) adopts a sparse mixture-of-experts architecture that activates only a subset of expert networks during inference, thereby maintaining strong performance with reduced computational cost. However, deterministic predictions usually cannot satisfy the requirement of decision-making. To address this limitation, Moirai (woo2024unified) trains a probabilistic model that captures a mixture of distributions. Building on language modeling techniques, Chronos (ansari2024chronos) discretizes time series through scaling and quantization, and models the resulting categorical distributions using cross-entropy loss. To enhance inference efficiency and forecasting accuracy, Chronos-Bolt replaces discrete tokenization with a patch-based input strategy and utilizes decoder representations to produce quantile forecasts over multiple future steps, achieving improved performance over its predecessor. These methods, however, lack inherent mechanisms to incorporate external contextual knowledge dynamically to facilitate zero-shot learning and suffer from limited interpretability.
@@
Retrieval-Augmented for Time Series Forecasting Although both LLMs and TSFMs have achieved strong performance in time series forecasting, they can still struggle in scenarios involving non-stationarity or distribution shifts. To address these challenges, Retrieval-Augmented Generation (RAG) techniques lewis2020retrieval offer a promising approach to incorporate external knowledge and enhance generalization and robustness. Recent works have explored the integration of retrieval mechanisms into time series forecasting. Among them, several approaches rely on fine-tuning to adapt the model to downstream tasks, including ReTime jing2022retrieval, RATD liu2024retrievaldiffusion, TimeRAG yang2024timerag, and RAFT han2025retrieval. ReTime jing2022retrieval proposes relational retrieval and content synthesis for retrieval-based time series forecasting. RATD liu2024retrievaldiffusion utilizes retrieved historical time series to guide the denoising process of diffusion models, enhancing forecasting performance. RAFT han2025retrieval integrates retrieval with a multi-resolution forecasting framework. These methods typically construct the retrieval database from the training set of the target task and require fine-tuning for model adaptation. Similarly, TimeRAG yang2024timerag incorporates retrieved sequences into the forecasting process by using a frozen LLM backbone, and introduces a trainable reprogramming layer to align time series and text modalities.
@@
RAF tire2024retrieval is a pioneering work that introduces a retrieval-augmented framework for zero-shot time series forecasting. It utilizes Chronos as the backbone model and constructs the augmented input by directly concatenating the processed retrieved context with the original time series input. However, this approach may face scalability and efficiency challenges when applied to large-scale retrieval databases. Moreover, RAG techniques for time series forecasting remain underexplored, particularly in terms of knowledge base construction, augmentation strategies, and their integration with TSFMs to facilitate zero-shot forecasting. To bridge these gaps, we propose TS-RAG, a retrieval-augmented framework specifically designed to enhance zero-shot forecasting capabilities of TSFMs.
@@ 3.1 Construction of the Retrieval Knowledge Base
TSFMs are typically pretrained on a multi-domain time series dataset to enhance their generalization capability. For TS-RAG, we adopt a similar strategy, i.e., construct a multi-domain dataset and specifically focus on learning the ARM augmentation module. We leverage the pretraining dataset of Chronos (ansari2024chronos), which utilizes TSMixup to randomly combine time series data points from various domains. This approach enhances data diversity by blending different patterns, thereby improving the model’s ability to generalize. Given the fact that the trainable parameters of TS-RAG are significantly less than those in the TSFM backbone and the retrieval encoder (more discussions are available in Appendix B.4), we uniformly sample a subset from the Chronos pretraining dataset to serve as the pretraining dataset for TS-RAG. Based on this subset, we can further draw a subset to construct the retrieval knowledge base for TS-RAG, which will be used in the inference (forecasting) stage.
@@ 3.2 Architecture of TS-RAG Framework
Within TS-RAG, a TSFM has three key components (tan_are_2024): an encoding layer, which may include normalization and embedding layers to preprocess and transform the input time series; a backbone, typically implemented as a transformer based model (i.e. GPT (radford2018improving), T5 (raffel2020exploring), Llama (touvron2023llama2) etc.) to extract temporal representations; and a projection layer, often implemented as a multi-layer perceptron (MLP), which maps the temporal representations from the backbone to the final prediction values.
@@ 4.1 Experimental Setup
Datasets and Retrieval Knowledge Base. For the pretraining dataset, we first uniformly sample 50 million data points from the Chronos pretraining dataset ansari2024chronos and further uniformly sample a subset of 5 million data points to construct the multi-domain retrieval knowledge base. To facilitate efficient indexing and retrieval, both the pretraining dataset and the retrieval knowledge base are segmented using a predefined context window. This process results in a total of 26 million pretraining data pairs and 2.8 million retrieval knowledge base pairs.
@@
Baselines. In practice, we use Chronos-Bolt, one of the state-of-the-art TSFMs, as the backbone of TS-RAG, as it achieves competitive performance in our evaluations. TS-RAG is designed to be compatible with any general TSFM. While our main experiments primarily use Chronos-Bolt as the backbone due to its strong empirical performance, we also verify TS-RAG’s effectiveness with MOMENT (goswami2024moment) in Appendix B.1. For comparison, we also report the zero-shot performance of other TSFMs, including TTM (ekambaram2024tiny), TimesFM (das2023decoder), Moirai (woo2024unified), Chronos (ansari2024chronos), Chronos-Bolt (ansari2024chronos), MOMENT (goswami2024moment), and Time-MoE shi2024timemoe.
@@ 4.3 Ablation Studies
We further conduct additional ablation studies to investigate the impact of retriever configurations, including the number of retrieved sequences, retriever encoder choice, and retrieval distance metrics. The results confirm that while TS-RAG benefits from a moderate number of retrieved sequences, it is robust to the choice of retriever encoder and retrieval distance metric. Detailed results and analysis are provided in Appendices B.3, B.4, and B.5.
@@ 4.4 Comparison with RAF
To further evaluate the performance of TS-RAG, we compare it with RAF (Retrieval Augmented Forecasting) tire2024retrieval, a recent retrieval-augmented method for zero-shot time series forecasting. As shown in Table 4, we conduct the comparison from two aspects: effectiveness and efficiency. More implementation details can be found in Appendix A.3.
@@ 5 Conclusion
Acknowledgements
The authors gratefully acknowledge the support from Morgan Stanley. Part of this work was conducted during Kanghui Ning’s internship at Ant Group.
@@ A.1 TS-RAG Implementation Details
During pretraining, all parameters of the TSFM backbone are frozen, and only the additional parameters introduced by TS-RAG are fine-tuned. The same forecasting loss as the backbone TSFM is adopted; specifically, when using Chronos-Bolt as the backbone, we apply the quantile regression loss following its original implementation. The number of retrieved sequences ($\text{top-}k$) is set to 10 by default; however, due to the flexible design of the ARM augmentation module, different values of $k$ can be explored. The model is trained using the AdamW optimizer with a learning rate of 0.0003 and a weight decay of 0.01. The batch size is set to 256, and training is conducted for 10,000 steps. To improve generalization, dropout is applied to certain layers with a dropout rate of 0.2. The training process was conducted on an NVIDIA A6000-48G GPU using TF32 precision. For efficient retrieval, FAISS is used to quickly identify the most relevant sequences from the retrieval knowledge base.
@@ A.2 TSFM Baseline Introduction
  • Chronos-Bolt (ansari2024chronos): Chronos-Bolt is a subsequent version of Chronos, which can handle patch-based and uses decoder representations to generate quantile forecasts across multiple future steps, improving forecast accuracy over Chronos.
  • MOMENT (goswami2024moment): MOMENT uses a masking modeling technique for zero-shot forecasting by appending a lookback series with a mask that matches the length of the forecast. It involves pretraining a Transformer encoder model univariately on the "Time Series Pile" datasets, which include a wide variety of time series data.
  • TTM (ekambaram2024tiny): TTM pre-trains a compact model based on the lightweight TSMixer architecture. It incorporates adaptive patching, diverse resolution sampling, and resolution prefix tuning to pretrain successfully on a small dataset.
  • Moirai (woo2024unified): Moirai pretrains the Transformer encoder on the “LOTSA” dataset, which includes 27B time points, by masking the forecast horizon of each target channel and performing mask reconstruction.
  • TimesFM (das2023decoder): TimesFM employs a decoder-style attention model and is pre-trained in a univariate manner on a large group of both real-world and synthetic datasets.
  • Chronos (ansari2024chronos): Chronos is a probabilistic time series foundation model.Chronos tokenizes the input time series in a quantized manner and processes these tokens using the T5 model (raffel2020exploring). Chronos is trained on an extensive corpus of collected and synthetic time series data and has great generalization ability.
  • Time-MOE (shi2024timemoe): Time-MoE consists of a family of decoder-only time series foundation models with a mixture-of-experts architecture, designed to operate in an auto-regressive manner, enabling universal forecasting with arbitrary prediction horizons and long context lengths.